if the area of a rectangle is 243 square m and the length is 27 m, what is the width of the rectangle
Found 2 solutions by vheroli, lenny460:
Answer by vheroli(126)

A= L * W
243 = 27 * W : divide both sides by 27 to get the value of width
243/27 = (27 * W)/27
9= W
Answer: 9m

Answer by lenny460(1073)

if the area of a rectangle is 243 square m and the length is 27 m, what is the width of the rectangle
Area of a rectangle = length * width
27 * w = 243
Divide each side by 27
(27 * w)/27 = 243/27
width = 9m

Answer: 9m